OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? −x + 2y = 9 4x + 5y = 9 −9 + 2y −9 − 2y 9 − 2y 9 + 2y

OpenStudy (reemii):

what is the first equation when you isolate x ?

OpenStudy (reemii):

isolate the x : -x = 9- 2y iff x = -9 + 2y.